# UQGNN: Uncertainty Quantification of Graph Neural Networks for Multivariate Spatiotemporal Prediction

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure 33rd ACM SIGSPATIAL International Conference on Advances in Geographic Information Systems; November 03–06, 2025; Minneapolis, MN, USAisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Data miningccs: Computing methodologies Machine learning approaches
@@ 5.1.1. Datasets
We evaluate UQGNN on four real-world heterogeneous urban datasets from Shenzhen, New York City (NYC), and Chicago. The Shenzhen dataset includes five mobility modes (bus, subway, taxi, ridesharing, bikesharing) across 491 regions, recorded hourly from January to June 2019. For NYC, the first dataset covers taxi (NYC, Taxi), bikesharing (NYC, ring), and subway (NYC, bway); the second includes (NYC, rime) and crash (NYC, rash) data, both recorded every 15 minutes in 67 Manhattan zones from January to June 2022. The Chicago dataset (taxi (Chicago, Taxi), bikesharing (Chicago, ring), and crime (Chicago, rime)) spans 77 regions with 15-minute records from September to November 2022.
@@ 5.1.2. Baselines
We compare our UQGNN with 12 state-of-the-art baselines, including 7 deterministic spatiotemporal prediction models, i.e., Spatial-Temporal Graph Convolution Network (STGCN) (Yu et al., 2017), Diffusion Convolutional Recurrent Neural Net (DCRNN) (Li et al., 2017), Graph WaveNet (GWNET) (Wu et al., 2019), Spectral Temporal Graph Neural Network (StemGNN) (Cao et al., 2020), Dynamic Spatial-Temporal Aware Graph Neural Network (DSTAGNN) (Lan et al., 2022), Adaptive Graph Convolutional Recurrent Network (AGCRN) (Bai et al., 2020), and SUMformer (Cheng et al., 2025), and 5 probabilistic spatiotemporal prediction models, i.e., TimeGrad (Rasul et al., 2021), Spatial-Temporal Zero-Inflated Negative Binomial Graph Neural Network (STZINB) (Zhuang et al., 2022), Deep Spatio-Temporal Uncertainty Quantification (DeepSTUQ) (Qian et al., 2023), DiffSTG (Wen et al., 2023), and CF-CNN (Huang et al., 2024). Detailed descriptions of these baselines are in Appendix B.1.
@@ 5.2. Overall Performance Comparison (RQ 1)
As shown in Figure 5, UQGNN’s deterministic predictions align most closely with the ground truth. In the probabilistic setting, the shadow around the curve represents the prediction interval, where a tighter shadow indicates more reliable predictions. It is observed that the green shadow is more compact while still covering most observations, indicating more reliable and precise estimates compared to baselines.
@@ 6.2. Uncertainty-aware Spatiotemporal Prediction
Nevertheless, most existing works on uncertainty-aware spatiotemporal prediction only consider a single urban phenomenon without capturing interactions between different urban phenomena. Our data-driven results show both opportunities and challenges to utilize data from one urban phenomenon to enhance the prediction of another phenomenon caused by their complicated interactions, which motivates us to design a new uncertainty-aware GNN for multivariate spatiotemporal prediction.
@@ 7. Conclusion
In this work, we design an uncertainty-aware graph neural network framework called UQGNN for multivariate spatiotemporal prediction. There are two key novel designs in UQGNN, i.e., an Interaction-aware Spatiotemporal Embedding Module with two innovative GNNs for capturing complicated spatiotemporal interactions of heterogeneous urban phenomena and a Multivariate Probabilistic Prediction Module for the prediction of both expected values and uncertainty. This framework can be adapted to various multivariate distributions based on different data characteristics. We conduct extensive experiments to show the performance of UQGNN on four real-world multivariate spatiotemporal datasets from Shenzhen, NYC, and Chicago. We compare UQGNN with 12 state-of-the-art spatiotemporal prediction models of 6 metrics The results show UQGNN effectively outperforms baselines and improves prediction accuracy and uncertainty quantification by 5%.
@@ B.2.2. Probabilistic Prediction Metrics
    (15) $\text{CRPS}(F,x)=\int_{\mathbb{R}}\left(F(z)-\mathbbm{1}(z\geq x)\right)^{2}\,\text{d}z$
    where $\mathbbm{1}(z\geq x)$ is the Heaviside step function (indicator function) which equals $1$ if $z\geq x$, and 0 otherwise. Smaller CRPS means better performance.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3748636.3762709isbn: 979-8-4007-2086-4/2025/11ccs: Information systems Data miningccs: Computing methodologies Machine learning approaches
@@ 5.1.1. Datasets
We evaluate UQGNN on four real-world heterogeneous urban datasets from Shenzhen, New York City (NYC), and Chicago. The Shenzhen dataset includes five mobility modes (bus, subway, taxi, ridesharing, bikesharing) across 491 regions, recorded hourly from January to June 2019. For NYC, the first dataset covers taxi (NYC, Taxi), bikesharing (NYC, ring), and subway (NYC, bway); the second includes crime (NYC, rime) and crash (NYC, rash) data, both recorded every 15 minutes in 67 Manhattan zones from January to June 2022. The Chicago dataset (taxi (Chicago, Taxi), bikesharing (Chicago, ring), and crime (Chicago, rime)) spans 77 regions with 15-minute records from September to November 2022.
@@ 5.1.2. Baselines
We compare our UQGNN with 12 state-of-the-art baselines, including 7 deterministic spatiotemporal prediction models (i.e., Spatial-Temporal Graph Convolution Network (STGCN) (Yu et al., 2017), Diffusion Convolutional Recurrent Neural Net (DCRNN) (Li et al., 2017), Graph WaveNet (GWNET) (Wu et al., 2019), Spectral Temporal Graph Neural Network (StemGNN) (Cao et al., 2020), Dynamic Spatial-Temporal Aware Graph Neural Network (DSTAGNN) (Lan et al., 2022), Adaptive Graph Convolutional Recurrent Network (AGCRN) (Bai et al., 2020), and SUMformer (Cheng et al., 2025)), and 5 probabilistic spatiotemporal prediction models (i.e., TimeGrad (Rasul et al., 2021), Spatial-Temporal Zero-Inflated Negative Binomial Graph Neural Network (STZINB) (Zhuang et al., 2022), Deep Spatio-Temporal Uncertainty Quantification (DeepSTUQ) (Qian et al., 2023), DiffSTG (Wen et al., 2023), and CF-CNN (Huang et al., 2024)). Detailed descriptions of these baselines are provided in Appendix B.1.
@@ 5.2. Overall Performance Comparison (RQ 1)
As shown in Figure 5, UQGNN’s deterministic predictions align most closely with the ground truth. In the probabilistic setting, the shadow around the curve represents the prediction interval, where a tighter shadow indicates more reliable predictions. It is observed that the green shadow of our UQGNN is more compact while still covering most observations, indicating more reliable and precise estimates compared to baselines.
@@ 6.2. Uncertainty-aware Spatiotemporal Prediction
Nevertheless, most existing works on uncertainty-aware spatiotemporal prediction only consider a single urban phenomenon without capturing interactions between different urban phenomena. Our data-driven results show both opportunities and challenges to utilize data from one urban phenomenon to enhance the prediction of another phenomenon caused by their complicated interactions, which motivates us to design new uncertainty-aware GNNs for multivariate spatiotemporal prediction.
@@ 7. Conclusion
In this paper, we propose an uncertainty-aware graph neural network framework called UQGNN for multivariate spatiotemporal prediction. There are two key novel designs in UQGNN, i.e., an Interaction-aware Spatiotemporal Embedding module with two innovative GNNs for capturing complicated spatiotemporal interactions of heterogeneous urban phenomena and a Multivariate Probabilistic Prediction module for the prediction of both expected values and uncertainty. This framework can be adapted to various multivariate distributions based on different data characteristics. We conduct extensive experiments by comparing UQGNN with 12 state-of-the-art baselines in terms of 6 metrics using four real-world multivariate spatiotemporal datasets from Shenzhen, NYC, and Chicago. The experimental results show that our UQGNN effectively outperforms the baselines by improving the prediction accuracy and uncertainty quantification by 5%.
@@ B.2.2. Probabilistic Prediction Metrics
    (15) $\text{CRPS}(F,x)=\int_{\mathbb{R}}\left(F(z)-\mathds{1}(z\geq x)\right)^{2}\,\text{d}z$
    where $\mathds{1}(z\geq x)$ is the Heaviside step function (indicator function) which equals $1$ if $z\geq x$, and 0 otherwise. Smaller CRPS means better performance.